# Gaussian Process Regression as a Sustainable Data-driven Background Estimate Method at the (HL)-LHC

### v3

@@ 1 Introduction
Gaussian Process Regression (GPR), on the other hand, is a non-parametric technique broadly used in machine learning and statistics [26]. A Gaussian Process is a collection of random variables such that any finite combination of them has a joint Gaussian distribution. Its potential usage in HEP experiments is discussed in ref. [27, 28] and several LHC results have applied this method to achieve various goals, such as template smoothing [29] or background estimation [18, 30]. In a GPR, the bin contents $y_{1}...y_{n}$ are described by a Gaussian PDF:
@@ 2 Test Dataset
The majority of published resonance searches consider smoothly falling backgrounds. To ensure no local features are introduced by the trigger selection, the low-mass regions are often not used, abandoning a sizeable amount of data. There have been attempts to include the whole dataset, such as a search for $b$-tagged resonances conducted in CMS [31]. In this search, the $m_{\mathrm{jj}}$ spectrum is divided into three regions, with each region fitted by a different function because there are no proper single functions that can describe the entire spectrum. One clear advantage of GPR is the ability to deal with complex spectra; hence, this work selects this analysis and uses its published data to construct the test dataset. A large set of pseudo-experiments is generated for the statistical tests, using the smooth background extracted from the results presented in ref. [31], with a uniform bin width of 15 GeV. Each pseudo-dataset is achieved by varying the event count in every bin independently according to a Poisson law. The corresponding integrated luminosity is therefore 36 $\text{fb}^{-1}$ roughly, hereinafter referred to as the “LHC scenario”. Figure 1 shows the seed template used to generate the pseudo-data and one example dataset. The signal injection tests mentioned in Section 5 are generated similarly after the signal events are added to the background.
The binning of the dataset impacts the performance, and the choice is made based on several confounding factors, such as the detector resolutions and computing cost. Given the computing complexity of GPR is $\mathcal{O}(n^{3})$ for $n$ data points [26], and the foreseeable enormous dataset expected at the (HL-)LHC, exploring GPR using binned data is arguably more practical. A comparative study of the binning choice can offer valuable insights to the community, although it is not studied in this work.
@@ 3.1 Data Pre-processing
The dataset used for resonance searches in the hadronic final states is often very sizeable. In this work, we use a binned dataset, so for each bin centre ($m_{\mathrm{jj}}$) there is a certain number of entries ($N_{\mathrm{events}}$), corresponding to the input data points ($x$) and the ones to be predicted by GPR ($y$), respectively. Its rapidly changing nature and the large mass coverage require a dedicated pre-processing step. Applying the logarithms of $x$ and $y$ can achieve both fast convergence and good performance. In the case of empty bins where $\log(y_{i})$ is not defined, zero padding is adopted. Figure 2 shows an example pre-processed background dataset and hypothetical Gaussian-shaped signal events. A Gaussian-shaped signal approximates a resonance with a given mass at the mean. The width of the resonance is parametrised by the ratio between the standard deviation and the mean. A benchmark value of 5% is chosen as it is a frequently considered scenario in resonance searches. The signal is well contained within an approximately 0.5 interval on the $x$-axis, motivating the hyperparameter choice made in Section 3.4.
@@ 3.2 GPR Model Setup
The kernel choice has a significant impact on the performance. As already mentioned in Section 1, there is a large series of kernels to choose from [26]. The goal of this work is not to find or construct the most optimal kernel, but to demonstrate the generality of a GPR-based method. Therefore, we use the multiplication of the two most common ones, an RBF and a constant kernel: $\mathrm{C}(c)\times\mathrm{RBF}(\ell)$. Previous works [27, 28] have shown this combination is a viable choice for HEP experiments. The only two parameters to be optimised are $c$ and $\ell$, and the bounds for $c$ and $\ell$ are hyperparameters defined a priori. The length scale of the RBF kernel, $\ell$, determines the smoothness of the model, and it has to be large enough in order not to incorporate the signal events. It is because a localised feature like a narrow signal will create bumpy features. The prior mean of $y_{i}$ has a diminishing impact on the posterior distribution, and is set to zero.
@@ 5 Sensitivity Study
Previously, we have identified a configuration that gives us acceptable background modelling precision in the background-only case. This section will demonstrate that this configuration retains good sensitivity to signals across the entire spectrum. When it comes to analyses using data-driven background estimate methods, we have to check carefully how sensitive the method is to the possible signal events. It is evaluated by a series of signal injection tests with Gaussian-shaped signal events. Four mass points are included to cover various locations in the spectrum. The 250 GeV point is below the trigger threshold, and the 300 GeV point is right at the plateau. These regions are often discarded in physics analyses. The 1400 GeV point is close to the end of the $m_{\mathrm{jj}}$ distribution, while the 800 GeV point is in the middle of the smoothly falling region where optimal sensitivity is expected when a traditional functional fit strategy is applied.
@@ 5.1 Sensitivity Evaluation
Figure 7 presents the chance of successfully reporting a $p$-value less than 0.1 in the correct mass region where the signal events are injected. It is noticed that when injecting a 300 GeV signal, there is a high chance of reporting a significant deviation at the wrong location. It is related to the residual bias near 400 GeV seen in Figure 5. Such mis-modelling effects should be taken into account as systematic uncertainties, or resolved via further tuning of the model parameters. Figure 12 in Section 6 observes a similar but predominating effect, which is mitigated by updated multiplication factors ($f_{i}$) as seen in Appendix A. Figure 8 shows random examples of the pseudo-experiments that have successfully reported a $p$-value below 0.1, with the BH intervals marked as well.
@@
Figure 9: An example of signal extraction test done for the 800 GeV Gaussian-shaped signal with a 5% width (left), and the summary of test results with different injected signal strengths (right), in the LHC scenario. The numbers of injected signal events are indicated by the green stars, while the blue crosses represent the medians of extracted signal. The width of the shaded bands corresponds to the density of a given number of extracted signal events.
@@ 6 Robustness in HL-LHC
  • Fit the 36.1 $\text{fb}^{-1}$ $m_{\mathrm{jj}}$ spectrum obtained by GPR using $f(x)=p_{0}(1-x)^{p_{1}}x^{p_{2}}$, where $x=m_{\mathrm{jj}}/6500$, starting from $m_{\mathrm{jj}}$ = 1000 GeV.
@@
Figure 15: An example of signal extraction test done for the 2000 GeV Gaussian-shaped signal with a 5% width (left), and the summary of test results with different injected signal strengths (right), in the HL-LHC scenario. The numbers of injected signal events are indicated by the green stars, while the blue crosses represent the medians of extracted signal. The width of the shaded bands corresponds to the density of a given number of extracted signal events.
The overall performance of the same GPR model decreases slightly in the HL-LHC scenario if the hyperparameters are not re-optimised (HL-LHC results using re-optimised hyperparameters are discussed in Appendix A). It is remarkable as the luminosity is increased by two orders of magnitude. The hyperparameters permit the users to tune the model with great flexibility, while the traditional functional fit method limits the users to a given function family. Due to the sizeable correlations between the free parameters, the fit performance of a function is saturated after reaching a certain number of free parameters. In contrast, the GPR-based background modelling technique can offer a solution that is valid through the entire lifetime of the LHC.
@@ Acknowledgments
Appendix A Impacts of the $L_{2}$ regularisation

### v4

@@ 1 Introduction
Gaussian Process Regression (GPR), on the other hand, is a non-parametric technique broadly used in machine learning and statistics [26]. A Gaussian Process is a collection of random variables such that any finite combination of them has a joint Gaussian distribution. Its potential usage in HEP experiments is discussed in ref. [27, 28] and several LHC results have applied this method to achieve various goals, such as template smoothing [29] or background estimation [18, 30]. In a GPR considering binned data, the bin contents $y_{1}...y_{n}$ are described by a Gaussian PDF:
@@ 2 Test Dataset
The majority of published resonance searches consider smoothly falling backgrounds. To ensure no local features are introduced by the trigger selection, the low-mass regions are often not used, abandoning a sizeable amount of data. There have been attempts to include the whole dataset, such as a search for $b$-tagged resonances conducted in CMS [31]. In this search, the $m_{\mathrm{jj}}$ spectrum is divided into three regions, with each region fitted by a different function because there are no proper single functions that can describe the entire spectrum. One clear advantage of GPR is the ability to deal with complex spectra; hence, this work selects this analysis and uses its published data to construct the test dataset. A large set of pseudo-experiments is generated for the statistical tests, using the smooth background extracted from the results presented in ref. [31], with a uniform bin width of 15 GeV. Each pseudo-dataset is achieved by varying the event count in every bin independently according to a Poisson law. The corresponding integrated luminosity is 36 $\text{fb}^{-1}$ roughly, hereinafter referred to as the “LHC scenario”. Figure 1 shows the seed template used to generate the pseudo-data and one example dataset. The signal injection tests mentioned in Section 5 are generated similarly after the signal events are added to the background.
The binning of the dataset impacts the performance, and the choice is made based on several confounding factors, such as the detector resolutions and computing cost. Given the computing complexity of GPR is $\mathcal{O}(n^{3})$ for $n$ data points [26], and the foreseeable enormous dataset expected at the (HL-)LHC, exploring GPR using binned data is arguably more practical than considering unbinned data. A comparative study of the binning choice can offer valuable insights to the community, although it is not studied in this work.
@@ 3.1 Data Pre-processing
The dataset used for resonance searches in the hadronic final states is often very sizeable. In this work, we use a binned dataset, so for each bin centre ($m_{\mathrm{jj}}$) there is a certain number of entries ($N_{\mathrm{events}}$), corresponding to the input data points ($x_{i}$) and the ones to be predicted by GPR ($y_{i}$), respectively. Its rapidly changing nature and the large mass coverage require a dedicated pre-processing step. Applying the logarithms of $x$ and $y$ can achieve both fast convergence and good performance. In the case of empty bins where $\log(y_{i})$ is not defined, zero padding is adopted. Figure 2 shows an example pre-processed background dataset and hypothetical Gaussian-shaped signal events. A Gaussian-shaped signal approximates a resonance with a given mass at the mean. The width of the resonance is parametrised by the ratio between the standard deviation and the mean. A benchmark value of 5% is chosen as it is a frequently considered scenario in resonance searches. The signal is well contained within an approximately 0.5 interval on the $x$-axis, motivating the hyperparameter choice made in Section 3.4.
@@ 3.2 GPR Model Setup
The kernel choice has a significant impact on the performance. As already mentioned in Section 1, there is a large series of kernels to choose from [26]. The goal of this work is not to find or construct the most optimal kernel, but to demonstrate the generality of a GPR-based method. Therefore, we use the multiplication of the two most common ones, an RBF and a constant kernel: $\mathrm{C}(c)\times\mathrm{RBF}(\ell)$. Previous works [27, 28] have shown this combination is a viable choice for HEP experiments. The only two parameters to be optimised are $c$ and $\ell$, and the bounds for $c$ and $\ell$ are hyperparameters defined a priori. The length scale of the RBF kernel, $\ell$, determines the smoothness of the model, and it has to be large enough in order not to incorporate the signal events. It is because a localised feature like a narrow signal will create bumpy features. The prior mean of $y_{i}$ ($u_{i}$) has a diminishing impact on the posterior distribution, and is set to zero.
@@ 5 Sensitivity Study
Previously, we have identified a configuration that gives us acceptable background modelling precision in the background-only case. This section will demonstrate that this configuration retains good sensitivity to signals across the entire spectrum. When it comes to analyses using data-driven background estimate methods, we have to check carefully how sensitive the method is to the possible signal events. It is evaluated by a series of signal injection tests with Gaussian-shaped signal events. Four mass points are included to cover various locations in the spectrum. The 250 GeV point is below the trigger threshold, and the 300 GeV point is right at the plateau. These regions are often discarded in physics analyses. The 1400 GeV point is close to the end of the $m_{\mathrm{jj}}$ distribution, while the 800 GeV point is in the middle of the smoothly falling region where optimal sensitivity is expected when a traditional functional fit strategy is applied. Appendix A discusses those various injection cases in further detail.
@@ 5.1 Sensitivity Evaluation
Figure 7 presents the chance of successfully reporting a $p$-value less than 0.1 in the correct mass region where the signal events are injected. It is noticed that when injecting a 300 GeV signal, there is a high chance of reporting a significant deviation at the wrong location. It is related to the residual bias near 400 GeV seen in Figure 5. Such mis-modelling effects should be taken into account as systematic uncertainties, or resolved via further tuning of the model parameters. Figure 12 in Section 6 observes a similar but predominating effect, which is mitigated by updated multiplication factors ($f_{i}$) as seen in Appendix B. Figure 8 shows random examples of the pseudo-experiments that have successfully reported a $p$-value below 0.1, with the BH intervals marked as well.
@@
Figure 9: An example of signal extraction test done for the 800 GeV Gaussian-shaped signal with a 5% width (left), and the summary of test results with different injected signal strengths (right), in the LHC scenario. The numbers of injected signal events are indicated by the green stars, while the blue dots represent the means of extracted signal. The width of the shaded bands corresponds to the density of a given number of extracted signal events. The red dotted lines are the one standard deviation boundaries.
@@ 6 Robustness in HL-LHC
  • Fit the 36.1 $\text{fb}^{-1}$ $m_{\mathrm{jj}}$ spectrum obtained by GPR using $f(x)=p_{0}(1-x)^{p_{1}}x^{p_{2}}$, where $x=m_{\mathrm{jj}}/6500$, starting from $m_{\mathrm{jj}}$ = 1000 GeV. $f(x)$ is the 3-parameter version of Function 1.
@@
Figure 15: An example of signal extraction test done for the 2000 GeV Gaussian-shaped signal with a 5% width (left), and the summary of test results with different injected signal strengths (right), in the HL-LHC scenario. The numbers of injected signal events are indicated by the green stars, while the blue dots represent the means of extracted signal. The width of the shaded bands corresponds to the density of a given number of extracted signal events. The red dotted lines are the one standard deviation boundaries.
The overall performance of the same GPR model decreases slightly in the HL-LHC scenario if the hyperparameters are not re-optimised (HL-LHC results using re-optimised hyperparameters are discussed in Appendix B). It is remarkable as the luminosity is increased by two orders of magnitude. The hyperparameters permit the users to tune the model with great flexibility, while the traditional functional fit method limits the users to a given function family. Due to the sizeable correlations between the free parameters, the fit performance of a function is saturated after reaching a certain number of free parameters. In contrast, the GPR-based background modelling technique can offer a solution that is valid through the entire lifetime of the LHC.
@@ Acknowledgments
Appendix A Signal Injected Spectra
The intrinsic difficulty to retain good sensitivity in the region below the smoothly falling part of the spectrum can be appreciated with examples such as Figure 16. When injecting a 5% Gaussian-shaped signal with a mass of 250 GeV, the injected spectrum has no clear localised structures. It is because this mass point is very close to the starting point of the fit, and the spectrum has a much steeper slope. Similarly, if a 300 GeV signal is considered, which is right at the plateau of the background, no obvious localised structures present neither. Those cases challenge the very assumption of the search strategy that the presence of a significant signal will create bumps. The ideal scenario is the 800 GeV signal, where the background on both sides follow the same trend. As the signal mass gradually approaches the end of the background distribution, the fit starts lacking constraints on the high mass side and suffering from large statistical fluctuations.
The above statements apply to both the functional fit method and the GPR-based method. Despite their entirely different underlying methodologies, they both rely solely on data. If the signal injected background spectrum is as smooth as the background-only one, neither method ought to achieve optimal sensitivity.
Figure 16: Examples of signal injected spectra for the 250, 300, 800 and 1400 GeV Gaussian-shaped signals with a 5% width. Both the background-only (solid black line), the 5-$\sigma$ injection (blue dashed line) and the 15-$\sigma$ injection (red dotted-dashed line) spectra are shown. The mass range is zoomed in to better visualise the region near the signal.
Appendix B Impacts of the $L_{2}$ regularisation